OpenStudy (anonymous):

multiply (8+√2)(4-3√2)

jimthompson5910 (jim_thompson5910):

(8+√2)(4-3√2) (8)(4)+(8)(-3√2)+(√2)(4)+(√2)(-3√2) 32-24√2+4√2-3*2 32-24√2+4√2-6 26-20√2 So \[\large (8+\sqrt{2})(4-3\sqrt{2})=26-20\sqrt{2}\]